### v1

@@ Abstract
Video Exposure Correction Quality Assessment, Vision-Language Guidance, Human Visual System
@@ 1 Introduction
All algorithms need a metric for improvement, where the quality assessment methods [21, 68, 76, 77, 24, 23, 79, 20, 78] play an important role. For VQA, there are mainly two branches: subjective VQA and objective ones. Subjective VQA is to assess the video by human, which is naturally more expensive and time-consuming. Objective VQA can be divided into Full-Reference (FR) VQA [2], Reduced-Reference (RR) VQA [50] and No-Reference (NR) VQA [43] contingent on the amount of required pristine video information. Due to the difficulty in obtaining reference videos, NR-VQA enjoys the most attention from researchers. In the early development stages of NR-VQA, researchers often evaluate video quality based on handcrafted features, such as structure, texture, and statistical features. Recently, owing to the potential in practical applications, deep-learning-based NR-VQA models with large language models have progressively dominated the VQA field. However, most existing VQA models are designed for general purposes instead of exposure correction. Few models specifically evaluate the quality of videos correcteds by VEC algorithms, which is possibly due to the lack of corresponding datasets.
@@
Subsequently, we propose a quality assessment model for inproper-exposed video enhancement based on Light-VQA, named Light-VQA+, whose capability is demonstrated in Table 1. Since brightness and noise have the most significant impact on VEC-VQA, we specifically extract features related to brightness, noise, and brightness consistency to enhance the model’s capabilities. While Light-VQA utilizes the traditional handcrafted algorithms to extract such feartures, Light-VQA+ borrows the strength of Large Language Models (LLM) [40], leading to a more accurate and efficient way for extracting such features. Besides, we still need the semantic features and motion features extracted from deep neural network when evaluating the quality of a video. The features mentioned above can be categorized into two aspects: spatial and temporal ones. To fuse these features, a cross-attention module [13] is deployed to combine the information from different sources. Also, it is worth noting that when a human watches a video, he/she does not focus evenly on the entire video. Specific video clips would receive more attention than others, which follows the Human Visual System (HVS) [4]. To better imitate the HVS, a trainable attention weight is then introduced when obtaining the final quality score of a video. Extensive experiments validate the effectiveness of our network design.
@@
    Benefiting from the developed dataset, we propose a Light-VQA-based quality assessment model: Light-VQA+, which extracts brightness and noise features through CLIP with the vision-language guidances provided by delicately designed prompts. Then it combines spatial-temporal information via cross-attention, followed with a quality regression to obtain a quality score. Finally, a set of weights are employed to make our model agree with the HVS.
@@
    The proposed Light-VQA+ surpasses Light-VQA on VEC-QA dataset as well as other public datasets. We envision that Light-VQA+ holds significant promise as a pivotal metric for the assessment as well as development of VEC algorithms.
@@ 2.4 NR-VQA Models
Lately, Large Language Models (LLM) have gained researchers’ attention for their incredible performance in multimodal learning. MaxVQA [65] combines CLIP, DOVER [66], and Fast-VQA to extract features from the input videos, making the process of VQA explainable. Q-Align [67] borrows the strength of LLaVA [29, 28, 30] as well as LoRA [17] to evaluate the quality of all image and video related contents. With well-designed prompts and powerful inference abilities, Q-Align outperforms many deep-learning-based methods in many downstream tasks.
While prior VQA models are designed for general UGC videos without exception, our model focuses on VEC quality assessment exclusively. We also utilize the large laguage model CLIP to enhance the assessment accuracy, stability and interpretability of our model.
@@ 3.2 Subjective Experiment
A high-quality dataset is a prerequisite for a well-performed model. To collect accurate annotations of the VEC-QA, we invite 21 experienced data labeling evaluators for a subjective experiment. Participants are tasked with rating video quality on a scale from 0 to 100, where a higher score indicates better quality. To ensure evaluators focus on perceptual quality rather than content, we guild the evaluators with a custom scoring process, details of which are shown in Algorithm 1. Participants are instructed to first rate the original video, followed by its corrected versions, with the option to repeatedly compare them with the original. This method, contrary to random shuffling, more effectively captures the perceptual differences due to exposure correction in the subjective quality assessments. We customize a scoring interface, which is demonstrated in Fig. 2.
@@ 4 Proposed Method
Benefiting from the constructed VEC-QA dataset as well as the success of Light-VQA [8], we further update the model into Light-VQA+, a multi-dimensional quality assessment model specialized in assessing exposure-corrected videos with vision-language guidance. This enhanced model explores the ability of large language models and cross-attention module, whose structure is depicted in Fig. 3. At the beginning, we divide the input video into 8 clips. Then, on each video clip, with the purpose of assessing them from a more comprehensive perspective, we extract both the spatial and temporal information for evaluation. To be specific, the Spatial Information (SI) is composed of deep-learning-based semantic features via Swin-Transformer [34] and CLIP [40]-captured brightness and noise features, while the Temporal Information (TI) consists of deep-learning-based motion features via SlowFast Model [10] and CLIP-captured brightness consistency features. These extracted features need a fusion procedure to create a quality-aware representation since they are generated from different perspectives. To this end, a cross-attention module [13] is applied to integrate them, followed by a two Fully-Connected (FC) layers that is designed for regressing these fused features into a unified video quality score for each clip. Last but not least, after obtaining a quality score for each video clip, we assign these video clips a set of trainable weights that are trained to be consistent with HVS [4]. Based on these weights, the weighted average of different video clips is calculated in the final score prediction.
@@ 4.1 Spatial Information
Figure 5: The structure of extracting the BNF and BCF from the third video clip. The “sub-videos” squared in red stand for the features that cover the third video clip. [Key: $V^{p}_{l,k}$: sub-videos; VF: Video Frames; BNF: Brightness & Noise Features; BCF: Brightness Consistency Features]
@@ 4.2 Temporal Information
To address this limitation, we implement a novel 4-level view strategy that focuses more comprehensively on the entire video rather than isolated video clips. We suppose the video contains $8n$ frames and employ a systematic approach to break down and reassemble the video into various “sub-videos” across different levels. For each level $l$, the video is divided into $2^{l-1}$ parts, with each part generating $2^{4-l}$ “sub-videos”. Each ”sub-video” captures frames with a stride of $2^{4-L}$, starting from the frame index $k$:
@@
where $l$ is the level index (1 to 4), $k\in\{1,...,2^{4-l}\}$, $t\in\{0,1,...,n-1\}$, and $p\in\{0,1,...,2^{l-1}-1\}$. $VF_{x}$ denotes the $x$-th frame of the entire video.
@@
4.3 Cross-Attention Fusion Module
After acquiring both the SI and TI, it becomes important to fuse them to achieve a more comprehensive representation of features, given that they originate from different perspectives. The cross-attention module [13], known for its effectiveness in integrating features from disparate sources, plays an important role in many tasks, such as in Stable Diffusion [42]. Consequently, in this paper, we employ the cross-attention module as our fusion module to amalgamate the SI and TI effectively. Specifically, we first normalize the dimensions of the SI and TI by two linear layers, denoted as $\mathcal{A}$ and $\mathcal{B}$, respectively. Following this, we apply a multi-head cross-attention module, denoted as $\mathcal{C}$. Specifically, the fusion process is conducted in two steps to underscore the equal importance of both the SI and TI. In the first step, we treat the normalized SI as both keys and values, with the normalized TI serving as the query. In the second step, we reverse these roles: the normalized TI is used as keys and values, while the normalized SI functions as the query. This symmetrical approach allows us to derive two distinct features, emphasizing the balanced integration of both types of information. Subsequently, we concatenate these two features and apply an additional linear layer, denoted as $\mathcal{F}$, to produce the final fused feature. The overall feature fusion module are represented as follows:
@@ 4.4 Quality Regression
Figure 6: The training attention weights from different models.
@@
where $Q_{i}$ denotes the quality score for the $i$-th video clip. Upon calculating the quality scores for each of the video clips, we introduce eight trainable parameters into our model to perform a weighted average calculation of these scores. These parameters allow the model to learn the relative importance of each video clip in determining the overall video quality, thus allocating more weight to those clips that have a greater impact. This approach is designed to align with the HVS, which naturally emphasizes more significant visual elements.
@@
where $Q$ represents the overall quality score of the video, and $w_{i}$ denotes the weight assigned to the $i$-th video clip’s quality score. To evaluate the effectiveness of this method, we conducted a series of experiments. Initially, identical weights were assigned to all eight video clips. This method was then applied across various models, and we tracked the convergence of their parameters. The results demonstrate that, irrespective of the model used, the final weights tend to be similar, as depicted in Figure 6. These results are consistent with the Human Visual System (HVS), indicating that our method is effective.
@@ 4.5 Loss Function
where $Q_{m}$, $\hat{Q}_{m}$ represents the ground truth and predicted MOS for the $m$-th video within a batch respectively, and $N$ denotes the batch size.
@@ 5.2 Performance Comparisons
Table 3: Experimental performance on our constructed VEC dataset along with its subset. Our proposed Light-VQA+ achieves the best performance. “HC”, “DL” and “LLM” denote Hand-Crafted-based, Deep-Learning-based and Large-Language-Models-based features respectively. The handcrafted models are inferior to deep-learning-based models, and deep-learning-based models are inferior to LLM-involved models. [Key: Best; Second Best]
VQA Model HC DL LLM LLVE-QA OEVR-QA VEC-QA
@@
Figure 8: The scatter plots of the predicted scores v.s. the MOS. The curves are obtained by a four-order polynomial nonlinear fitting. For a model, there are 3 scatter plots, standing for the testing results on VEC-QA, LLVE-QA, OEVR-QA respectively. It is evident that the predicted scores of our proposed VQA bear the closest resemblance to the MOS.
@@ 5.3 Cross-Dataset Validation
Table 4: Cross dataset validation on different VQA datasets with models trained on our VEC-QA dataset. Our proposed Light-VQA+ achieves the best performance. “HC”, “DL” and “LLM” denote three types of leveraged features, Hand-Crafted based, Deep-Learning based and Large-Language-Models based repectively. The handcrafted models are inferior to deep-learning-based models, and deep-learning-based models are inferior to LLM-involved models. [Key: Best; Second Best]
VQA Model HC DL LLM Mixed-A Mixed-B Mixed-C
@@
To evaluate the cross-dataset performance of the model, we conducted experiments using mixed subsets of videos with improper exposure from three sources: VDPVE [11], KoNViD_1k [16], and Live-VQC [48]. After selecting videos from these datasets, we combined them and randomly divided them into three subsets, designated as Mixed-A, Mixed-B, and Mixed-C, each containing 600 videos. We then directly applied models pre-trained on the VEC-QA dataset to test these new subsets, facilitating an efficient assessment process. The overall experimental results are presented in Table 4.
It is important to note that the VEC-QA dataset is specifically designed to include videos with improper exposure as well as their corrected counterparts, making it suitable for training exposure-related quality assessment models. However, the test subsets are not originally intended for exposure-focused tasks. This mismatch in dataset design has led to less effective performance of the quality-aware representations learned from the VEC-QA dataset, resulting in a performance decline across all tested methods. Despite this general downturn, our proposed Light-VQA+ method still managed to outperform the other seven VQA methods, showcasing its robust generalization ability in assessing the quality of exposure-corrected videos.
@@ 5.4 Ablation Studies
Table 5: Experimental performance of ablation studies on our constructed VEC dataset along with its subset. [Key: Best; Second Best]
Model BN & BC Fusion Attn-W LLVE-QA OEVR-QA VEC-QA
@@
3 LLM (Local) MLP 0.9231 0.9261 8.0513 0.7118 0.7260 8.7087 0.9010 0.9343 8.4690
4 LLM (Global) MLP 0.9231 0.9260 8.0588 0.7098 0.7316 8.6338 0.9021 0.7378 0.9350
5 LLM (Both) MLP 0.9317 0.9324 7.7131 0.7061 0.7264 8.7039 0.9010 0.9367 8.3187
6 LLM (Both) Cross-Attn 0.9290 0.9310 7.7910 0.7215 0.7457 8.4386 0.9056 0.9384 8.2094
7 LLM (Both) Cross-Attn 0.9428 0.9450 6.9803 0.7462 0.7772 7.9690 0.9177 0.9480 7.5597
In this subsection, we conduct a series of ablation experiments to evaluate the individual contributions of different modules within Light-VQA+. The results of these ablation studies are presented in Table 5. Model 1 removes the BNF and BCF extracted via CLIP, the cross-attention module [13] and the weights for aligning the HVS. Based on Model 1, Model 2 utilizes the handcrafted methods to extract the BNF and BCF. Model 3 utilizes the local features extracted by CLIP, which are the local BNF in SI and the Lv4 BCF in TI. Opposing to Model 3, Model 4 exploits the global features (i.e., the global BNF and the Lv1 to Lv3 BCF). Model 5 removes the cross-attention module and applies a MLP [53] instead, while Model 6 uses a simple average to replace the parameters that align with the HVS. Model 7 is the complete model we propose, in which we fuse all the spatial and temporal information, and obtain the best results. Based on the results of the 7 models, we can analyze the contribution of each modules in Light-VQA+.

### v2

@@ Abstract
Exposure Correction, Video Quality Assessment, Vision-Language Guidance, Human Visual System
@@ 1 Introduction
All algorithms need a metric for improvement, where the quality assessment methods [35, 36, 37, 38, 39, 40, 41, 42, 43] play an important role. For VQA, there are mainly two branches: subjective VQA and objective ones. Subjective VQA is to assess the video by human, which is naturally more expensive and time-consuming. Objective VQA can be divided into Full-Reference (FR) VQA [44], Reduced-Reference (RR) VQA [45] and No-Reference (NR) VQA [46] contingent on the amount of required pristine video information. Due to the difficulty in obtaining reference videos, NR-VQA enjoys the most attention from researchers. In the early development stages of NR-VQA, researchers often evaluate video quality based on handcrafted features, such as structure, texture, and statistical features. Recently, owing to the potential in practical applications, deep-learning-based NR-VQA models with multimodal large language models have progressively dominated the VQA field. However, most existing VQA models are designed for general purposes instead of exposure correction. Few models specifically evaluate the quality of videos correcteds by VEC algorithms, which is possibly due to the lack of corresponding datasets.
@@
Subsequently, we propose a quality assessment model for inproper-exposed video enhancement based on Light-VQA, named Light-VQA+, whose capability is demonstrated in Table 1. Since brightness and noise have the most significant impact on VEC-VQA, we specifically extract features related to brightness, noise, and brightness consistency to enhance the model’s capabilities. While Light-VQA utilizes the traditional handcrafted algorithms to extract such feartures, Light-VQA+ borrows the strength of Multimodal Large Language Models (MLLM) [47], leading to a more accurate and efficient way for extracting such features. Besides, we still need the semantic features and motion features extracted from deep neural network when evaluating the quality of a video. The features mentioned above can be categorized into two aspects: spatial and temporal ones. To fuse these features, a cross-attention module [48] is deployed to combine the information from different sources. Also, it is worth noting that when a human watches a video, he/she does not focus evenly on the entire video. Specific video clips would receive more attention than others, which follows the Human Visual System (HVS) [49]. To better imitate the HVS, a trainable attention weights is then introduced when obtaining the final quality score of a video. Extensive experiments validate the effectiveness of our network design.
@@
    Benefiting from the developed dataset, we propose a Light-VQA-based quality assessment model: Light-VQA+, which extracts brightness and noise features through CLIP with vision-language guidances provided by delicately designed prompts. Then it combines spatial-temporal information via cross-attention, followed with a quality regression to obtain the quality score for each video clip. Finally, a set of trainable weights are employed on all video clips to obtain the final score, making the assessment align with the HVS.
@@
    The proposed Light-VQA+ surpasses Light-VQA on VEC-QA dataset as well as other public datasets. We envision that Light-VQA+ holds significant promise as a pivotal metric for the assessment as well as the development of VEC algorithms.
@@ 2.4 NR-VQA Models
Lately, Multimodal Large Language Models (MLLM) have gained researchers’ attention for their incredible performance in multimodal learning. MaxVQA [3] combines CLIP, DOVER [71], and Fast-VQA to extract features from the input videos, making the process of VQA explainable. Q-Align [4] borrows the strength of LLaVA [72, 73, 74] as well as LoRA [75] to evaluate the quality of all image and video related contents. With well-designed prompts and powerful inference abilities, Q-Align outperforms many deep-learning-based methods in many downstream tasks.
While prior VQA models are designed for general UGC videos without exception, our model focuses on VEC quality assessment exclusively. We also utilize the multimodal large laguage model CLIP to enhance the assessment accuracy, stability and interpretability of our model.
@@ 3.2 Subjective Experiment
A high-quality dataset is a prerequisite for a well-performed model. To collect accurate annotations of the VEC-QA, we invite 21 experienced data labeling evaluators for a subjective experiment. Participants are tasked with rating video quality on a scale from 0 to 100, where a higher score indicates better quality. To ensure evaluators focus on perceptual quality rather than content, we guild the evaluators with a custom scoring process, details of which are shown in Algorithm 1. Participants are first required to evaluate the original video and then assess its corrected counterparts placed aside in a random order. Using this random order rather than a fixed one can more accurately reflect the perceptual differences attributed to exposure correction in the subjective experiment. To achieve this, we have developed a customized scoring interface shown in Fig. 2.
@@ 4 Proposed Method
Benefiting from the constructed VEC-QA dataset as well as the success of Light-VQA [5], we further update the model into Light-VQA+, a multi-dimensional quality assessment model specialized in assessing exposure-corrected videos with vision-language guidance. This enhanced model explores the ability of multimodal large language models, whose structure is depicted in Fig. 3. Among the open MLLMs with the ability of image-text interaction, CLIP and LLaVA [72, 73, 74] are the top performers to date. LLaVA employs the image-text pairing capabilities of CLIP, enhanced by integration with the strengths of LLaMA [76] in language processing. However, its capacity for image processing remains identical to that of CLIP. Given that our model does not require natural language generation, we choose to directly employ CLIP instead of LLaVA to conserve computational resources. At the beginning, we divide the input video into 8 clips. Then, on each video clip, with the purpose of assessing them from a more comprehensive perspective, we extract both the spatial and temporal information for evaluation. To be specific, the Spatial Information (SI) is composed of deep-learning-based semantic features via Swin-Transformer [69] and CLIP [47]-captured brightness and noise features, while the Temporal Information (TI) consists of deep-learning-based motion features via SlowFast Model [77] and CLIP-captured brightness consistency features. These extracted features need a fusion procedure to create a quality-aware representation since they are generated from different perspectives. To this end, a cross-attention module [48] is applied to integrate them, followed by a two Fully-Connected (FC) layers that is designed for regressing these fused features into a unified video quality score for each clip. Last but not least, after obtaining a quality score for each video clip, we assign these video clips a set of trainable weights that are autonomously learned to be consistent with HVS [49]. Based on HVS weights, the weighted average of different video clips is calculated to predict the final assessment score.
@@ 4.1 Spatial Information
Figure 5: The structure of extracting the BNF and BCF from the third video clip. The “sub-videos” $V^{p}_{l,k}$ squared in red stand for the features that cover the third video clip. [Key: VF: Video Frames; BNF: Brightness & Noise Features; BCF: Brightness Consistency Features]
@@ 4.2 Temporal Information
To address this limitation, we implement a novel 4-level view strategy that focuses more comprehensively on the entire video rather than isolated video clips. We suppose the video contains $8n$ frames and employ a systematic approach to break down and reassemble the video into various “sub-videos” across different levels. For each level $l$, the video is divided into $2^{l-1}$ parts, with each part generating $2^{4-l}$ “sub-videos”. Each “sub-video” captures frames with a stride of $2^{4-L}$, starting from the frame index $k$:
@@
where $l$ is the level index (1 to 4), $k\in\{1,2,...,2^{4-l}\}$, $t\in\{0,1,...,n-1\}$, and $p\in\{0,1,...,2^{l-1}-1\}$. $VF_{x}$ denotes the $x$-th frame of the entire video.
@@
4.3 Feature Fusion Module
After acquiring both SI and TI, it becomes important to fuse them to achieve a more comprehensive representation of features, given that they originate from different perspectives. The cross-attention module [48], known for its effectiveness in integrating features from disparate sources, plays an important role in many tasks, such as in Stable Diffusion [78]. Consequently, in this paper, we employ the cross-attention module as our fusion module to incorporate the SI with TI effectively. We first normalize the dimensions of the SI and TI by two linear layers, denoted as $\mathcal{A}$ and $\mathcal{B}$, respectively. Following this, we apply a multi-head cross-attention module, denoted as $\mathcal{C}$. Specifically, the fusion process is conducted in two steps to underscore the equal importance of both SI and TI. In the first step, we treat the normalized SI as both keys and values, with the normalized TI serving as the query. In the second step, we reverse these roles: the normalized TI is used as keys and values, while the normalized SI functions as the query. This symmetrical approach allows us to derive two distinct features, emphasizing the balanced integration of both types of information. Subsequently, we concatenate these two features and apply an additional linear layer, denoted as $\mathcal{F}$, to produce the final fused feature. The overall feature fusion module are represented as follows:
@@ 4.4 Quality Regression
Figure 6: The HVS weights under different models.
@@
where $Q_{i}$ denotes the quality score for the $i$-th video clip. Upon calculating the quality scores for each of the video clips, we introduce eight trainable parameters into our model to perform a weighted average calculation of these scores. These parameters allow the model to learn the relative importance of each video clip in determining the overall video quality, thus allocating more weight to those clips that have a greater impact.
@@
where $Q$ represents the overall quality score of the video, and $w_{i}$ denotes the weight assigned to the $i$-th video clip’s quality score.
To evaluate the effectiveness of this method, we conducted a series of experiments. Initially, identical HVS weights are assigned to all eight video clips. This method was then applied across various models, and we tracked the convergence of their parameters. The results demonstrate that, irrespective of the model used, the final HVS weights tend to be similar, as depicted in Figure 6. These results are consistent with the HVS that human perception usually pays more attention to the beginning of a video, confirming the effectiveness of our method.
(a) MOS
(b) Brightness
(c) Contrast
(d) Colorfulness
@@ 4.5 Loss Function
where $Q_{m}$, $\hat{Q}_{m}$ represent the ground-truth and the predicted MOS of the $m$-th video within a batch, and $N$ denotes the batch size.
@@ 5.2 Performance Comparisons
Table 3: Experimental performance on our constructed VEC dataset along with its subset. Our proposed Light-VQA+ achieves the best performance. “HC”, “DL” and “MLLM” denote Hand-Crafted-based, Deep-Learning-based and Large-Language-Models-based features respectively. The handcrafted models are inferior to deep-learning-based models, and deep-learning-based models are inferior to MLLM-involved models. [Key: Best; Second Best]
VQA Model HC DL MLLM LLVE-QA OEVR-QA VEC-QA
@@
Figure 8: The scatter plots of the predicted scores v.s. the MOS. The curves are obtained by a four-order polynomial nonlinear fitting. For a model, there are 3 scatter plots, standing for the testing results on VEC-QA, LLVE-QA, and OEVR-QA respectively. It is evident that the predicted scores of our proposed VQA bear the closest resemblance to the MOS.
@@ 5.3 Cross-Dataset Validation
Table 4: Cross-dataset validation on different VQA datasets. Our proposed Light-VQA+ achieves the best performance. “HC”, “DL” and “MLLM” denote Hand-Crafted-based, Deep-Learning-based and Large-Language-Models-based features respectively. The handcrafted models are inferior to deep-learning-based models, and deep-learning-based models are inferior to MLLM-involved models. [Key: Best; Second Best]
VQA Model HC DL MLLM Mixed-A Mixed-B Mixed-C
@@
To evaluate the cross-dataset performance of the model, we conduct experiments using mixed subsets of videos with improper exposure from three sources: VDPVE [57], KoNViD_1k [54], and Live-VQC [51]. After selecting videos from these datasets, we combine them and divided them into three subsets, designated as Mixed-A, Mixed-B, and Mixed-C from over-exposed, low-light, and merged perspectives, each containing 600 videos. We then directly apply models pre-trained on the VEC-QA dataset to test these new subsets, facilitating an efficient assessment process. The overall experimental results are presented in Tab. 4.
It is important to note that the VEC-QA dataset is specifically designed to include videos with improper exposure as well as their corrected counterparts, making it suitable for training exposure-related quality assessment models. However, the test subsets are not originally intended for this task. This mismatch in dataset design has led to less effective performance of the quality-aware representations learned from the VEC-QA dataset, resulting in a performance decline across all tested methods. Despite this general downturn, our proposed Light-VQA+ method still manages to outperform the other seven VQA methods, showcasing its robust generalization ability in assessing the quality of exposure-corrected videos.
@@ 5.4 Ablation Studies
Table 5: Experimental performance of ablation studies on our constructed VEC dataset along with its subset. [Key: Best; Second Best; HVS-W: HVS weights]
Model BN & BC Fusion HVS-W LLVE-QA OEVR-QA VEC-QA
@@
3 MLLM (Local) MLP 0.9231 0.9261 8.0513 0.7118 0.7260 8.7087 0.9010 0.9343 8.4690
4 MLLM (Global) MLP 0.9231 0.9260 8.0588 0.7098 0.7316 8.6338 0.9021 0.7378 0.9350
5 MLLM (Both) MLP 0.9317 0.9324 7.7131 0.7061 0.7264 8.7039 0.9010 0.9367 8.3187
6 MLLM (Both) Cross-Attn 0.9290 0.9310 7.7910 0.7215 0.7457 8.4386 0.9056 0.9384 8.2094
Light-VQA+ MLLM (Both) Cross-Attn 0.9428 0.9450 6.9803 0.7462 0.7772 7.9690 0.9177 0.9480 7.5597
In this subsection, we conduct a series of ablation studies to evaluate the individual contributions of different modules within Light-VQA+. The results of these ablation studies are presented in Tab. 5. Model 1 removes the BNF and BCF extracted via CLIP, the cross-attention module [48], and the HVS weights. Based on Model 1, Model 2 utilizes the handcrafted methods to extract the BNF and BCF. Model 3 utilizes the local features extracted by CLIP, which are the local BNF in SI and the Lv4 BCF in TI. Opposing to Model 3, Model 4 exploits the global features (i.e., the global BNF and the BCF from Lv1 to Lv3). Model 5 removes the cross-attention module and applies a MLP [81] instead, while Model 6 uses a simple average to replace the parameters that align with the HVS. The last one, Light-VQA+, is the complete model we propose, in which we fuse all the spatial and temporal information, and obtain the best results. Based on the results of the 7 models, we can analyze the contribution of each modules in Light-VQA+.